OpenStudy (anonymous):

5^2-7^2 -------- 5-7

OpenStudy (anonymous):

@mtrejo \[\frac{ (5^{2})(7^{2}) }{ 5-7 }\]\[\frac{ 25-49 }{ -2 }\]\[\frac{ -24 }{ -2 }\]Answer is:\[12\]